# ETGL-DDPG: A Deep Deterministic Policy Gradient Algorithm for Sparse Reward Continuous Control

### v1

@@ 1 Introduction
Deep deterministic policy gradient (DDPG) (Lillicrap et al., 2015) is one of the representative algorithms for reinforcement learning (RL) (Sutton & Barto, 2018), alongside other prominent approaches (Haarnoja et al., 2018; Fujimoto et al., 2018; Andrychowicz et al., 2017). The method has been extensively used for continuous control environments with dense reward signals (Duan et al., 2016). However, its performance degrades significantly when the reward signals are sparse and are only observed upon reaching the goal (Matheron et al., 2019).
@@
In this paper, we enhance DDPG (Lillicrap et al., 2015) to address all three aforementioned problems. Our first contribution is $\epsilon t$-greedy, a new temporally version of $\epsilon$-greedy that utilizes a light-weight search procedure, similar to Laud (2004), to enable more directional exploration based on the agent’s previous experience data. We show that similar to $\epsilon z$-greedy, $\epsilon t$-greedy has polynomial sample complexity in related parameters of the MDP. Our second contribution is a new goal-conditioned dual replay buffer (GDRB), that uses two replay buffers along with an adaptive sampling strategy to differentiate goal-reached and goal-not-reached experience data. These two buffers differ in retention policy, size, and the transitions they store. Our third enhancement is to replace the one-step update in DDPG with the longest $n$-step return for all transitions in an episode. Figure 1 illustrates the innovations of ETGL-DDPG. In Section 4, we evaluate the performance of ETGL-DDPG through extensive experiments on 2D and 3D continuous control benchmarks. We show that each of the three strategies individually improves the performance of DDPG. Furthermore, ETGL-DDPG outperforms current state-of-the-art methods across all tested environments.
@@ 3.1 $\epsilon t$-Greedy: Exploration with search
To execute the search process, the agent requires access to the environment’s transition function $\mathcal{T}$ of the corresponding MDP. This is used to generate new nodes within the search tree. However, since our exploration strategy is built on DDPG, the model-free algorithm, the transition function $\mathcal{T}$ is not known. Instead, the agent utilizes its replay buffer to advance the search. We briefly discuss the impact of having access to $\mathcal{T}$ on the exploration process in Supplementary Material A.2. We also assume that the agent has a SimHash function $\phi$, which discretizes the large continuous environment. For each state $s$, $n(\phi(s))$ serves as an estimate of the number of visits to a neighbourhood of $s$ throughout the entire learning process.
@@
A policy $\pi$ is called $\delta$-optimal if it satisfies $V^{\pi^{*}}(s)-V^{\pi}(s)\leq\epsilon$, for all $s\in S$, where $\epsilon>0$.
@@
For simplicity, when we say an algorithm $\mathcal{A}$ has polynomial sample complexity, we imply that $\mathcal{A}$ is PAC-MDP. The work by Liu & Brunskill (2018) establishes polynomial sample complexity for a uniformly random exploration by bounding the covering length defined in Definition 1. Using this, and considering a limited tree budget $N$, we show that $\epsilon t$-greedy is PAC-MDP. Let’s denote the search tree by $\mathcal{X}$, and the distribution over the generated options in $\mathcal{X}$ as $\mathcal{P}_{\omega}$. The following Theorem provides a lower bound on option sampling in tree $\mathcal{X}$ under certain condition.
@@ 3.3 Using Longest $n$-step return
$Q(s_{t},a_{t})=\begin{cases}\sum^{T-t}_{k=0}\gamma^{k}r_{t+k},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{$s_{T}$ is a goal state}\\ \\ \sum^{T-t-1}_{k=0}\gamma^{k}r_{t+k}+\gamma^{T-t}Q(s_{T},a_{T}),\ \ \ \ \ \text{otherwise}\end{cases}$ (6)
@@ 4.2 Environment Coverage through Exploration
We now examine how effective $\epsilon t$-greedy is in covering the environment. To do so, we discretize the navigation environments into small cells. A cell is considered visited if the agent encounters a sufficient number of distinct states within it, and the overall environment coverage is quantified as the fraction of visited cells. Figure 4 presents a comparison of environment coverage across different exploration strategies. All strategies except DOIE, which uses Radial Basis Function Deep Q-Network (RBFDQN) (Asadi et al., 2021), use DDPG as their underlying algorithm. RBFDQN is an enhanced DQN variant that incorporates Radial Basis Functions (RBF) to achieve more accurate Q-value approximations in continuous environments. In Wall-maze, $\epsilon t$-greedy is the only method capable of fully covering the environment, while DOIE achieves 90% coverage. $\epsilon z$-greedy covers approximately half of the environment, whereas the remaining methods manage to explore only around 30%. In U-maze, all strategies are successful, covering 80% or more of the environment. Even so, both $\epsilon t$-greedy and DOIE reach full coverage faster than other methods. In Point-push, none of the methods can fully cover the environment. However, $\epsilon t$-greedy still outperforms all baselines, and among the baselines, DOIE explores more than the others. We also investigate the distribution of final states reached in the episodes to determine the order in which the agent visits different regions of the environment (see Appendix A.5).
The tree budget $N$ upper bounds the option length of $\epsilon t$-greedy due to the fact that the longest path between nodes in the tree is shorter or equal to the number of nodes in the tree. This is analogous to the role of $N$ in $\epsilon z$-greedy, where a uniform distribution $z(n)=\mathbbm{1}_{n\leq N}/N$ is used. To evaluate both methods, we assess environment coverage under varying budget sizes, calculating the coverage after 1 million training frames. Table 1 shows the results: $\epsilon t$-greedy consistently achieves greater coverage than $\epsilon z$-greedy across all environments and budget sizes. Additionally, $\epsilon t$-greedy demonstrates improved the coverage as the budget increases. In contrast, increasing the budget for $\epsilon z$-greedy does not consistently improve coverage and can even decrease it in some cases. This highlights the advantages of directed exploration over undirected methods, particularly in complex environments with numerous obstacles, such as Wall-maze.
@@ 4.3 Effectiveness of Each New Component in ETGL-DDPG
We evaluated the performance of ETGL-DDPG, and now we assess the impact of each component on DDPG separately. Figure 5 presents the results for all environments. $\epsilon t$-greedy demonstrates the most improvement across all environments and is the only method that enhances the performance of DDPG in the Wall-maze, highlighting the critical role of our exploration strategy. GDRB shows a positive impact on DDPG performance in all environments, except for soccer, where DDPG alone outperforms all baselines. Additionally, we replaced reservoir sampling with FIFO as the retention policy in GDRB and observed similar results. The longest n-step return has a positive effect only in U-maze and press-button tasks, while it negatively impacts performance in soccer and Point-push. We attribute this to the inherently high variance of multi-step updates. A comparison of Figures 3 and 5 across all environments shows that ETGL-DDPG consistently outperforms the use of each component individually, supporting the effectiveness of their combination.
budget $N$
$\epsilon z$-greedy $\epsilon t$-greedy
Wall-maze U-maze Point-push Wall-maze U-maze Point-push
5 0.36 0.55 0.36 0.76 0.94 0.40
10 0.38 0.91 0.38 0.97 0.91 0.41
15 0.34 0.85 0.39 0.65 0.94 0.42
20 0.30 0.84 0.40 0.83 0.94 0.48
25 0.28 0.86 0.40 1 0.95 0.47
30 0.27 0.83 0.39 1 0.97 0.51
35 0.25 0.82 0.40 1 0.95 0.53
40 0.24 0.82 0.40 1 0.97 0.55
45 0.22 0.85 0.41 1 0.96 0.64
50 0.22 0.79 0.40 1 0.97 0.73
@@ 5 Related work
Exploration. Intrinsic motivation methods  (Burda et al., 2018; Pathak et al., 2017; Ostrovski et al., 2017; Tang et al., 2017) provide a reward bonus for unexplored areas of the state space. These methods make the reward function non-stationary, which breaks the Markov assumption of MDP. Decoupled RL algorithms (Schäfer et al., 2021; Badia et al., 2019) resolve the non-stationarity of the reward function by training two separate policies for exploration and exploitation. However, such methods require double the computation cost. Colas et al. (2018) use a policy search process to generate diverse data for training of DDPG. Liu et al. (2018) introduce a competition-based exploration method where two agents (A and B) compete with each other. Agent A is penalized for visiting states visited by B, while B is rewarded for visiting states discovered by A. Plappert et al. (2018) directly inject noise into the policy’s parameter space instead of the action space. Eysenbach et al. (2019) build a graph using states in the replay buffer, allowing the agent to navigate distant regions of the environment by applying Dijkstra’s algorithm. Lobel et al. (2022) present Deep Optimistic Initialization for Exploration (DOIE), which improves exploration in continuous control tasks by maintaining optimism in state-action value estimates. Lobel et al. (2023) demonstrate that DOIE can estimate visit counts by averaging samples from the Rademacher distribution instead of using density models. Dey et al. (2024) present COIN, a continual optimistic initialization strategy that extends DOIE to stochastic and non-stationary environments.
@@
Experience Replay Buffer and Reward Propagation. Rather than uniformly sampling from the buffer, Prioritized Experience Replay (PER) (Schaul et al., 2015) prioritizes transitions in the buffer based on reward, recency, or TD error at the expense of $O(\log N)$ per sample, where $N$ is the buffer size. CER (Zhang & Sutton, 2017) includes the last transition from the buffer to each sampled batch with $O(1)$ complexity. Zhang et al. (2022) learn a conservative value regularizer only from the observed transitions in the replay buffer to improve the sample efficiency of DQN. Pan et al. (2022) theoretically show why PER has a better convergence rate than uniform sampling policy when minimizing mean squared error. Furthermore, Pan et al. (2022) identify two limitations of PER: outdated priorities and insufficient coverage of the state space. Reward shaping (Laud, 2004; Hu et al., 2020) creates artificial intermediate rewards to facilitate reward propagation. However, designing appropriate intermediate rewards is hard and often problem-specific. Trott et al. (2019) address this issue by introducing  self-balancing reward shaping in the context of on-policy learning. To extract more information from an unsuccessful episode, Andrychowicz et al. (2017) introduce imaginary goals. An imaginary goal for state $s$ is a state that is encountered later in the episode. Devidze et al. (2024) introduce a novel reward informativeness criterion that adaptively designs interpretable reward functions based on an agent’s current policy in sparse-reward tasks.

### v2

@@ 1 Introduction
Deep deterministic policy gradient (DDPG) (Lillicrap et al., 2015) is one of the representative algorithms for reinforcement learning (RL) (Sutton & Barto, 2018), alongside other prominent approaches (Haarnoja et al., 2018; Fujimoto et al., 2018; Andrychowicz et al., 2017). In recent years, DDPG has served as the backbone algorithm for introducing novel ideas in robotics and RL Barth-Maron et al. (2018); Pan et al. (2020); Liu et al. (2023); Wang et al. (2023); Tiapkin et al. (2024). While DDPG demonstrates strong performance in continuous control tasks with dense reward signals (Duan et al., 2016; Kiran et al., 2021), its effectiveness diminishes significantly in sparse-reward settings where rewards are only observed upon reaching the goal (Matheron et al., 2019; Luck et al., 2019).
@@
In this paper, we enhance DDPG (Lillicrap et al., 2015) to address all three aforementioned problems. Our first contribution is $\epsilon t$-greedy, a new temporally version of $\epsilon$-greedy that utilizes a light-weight search procedure, similar to Laud (2004), to enable more directional exploration based on the agent’s previous experience data. We show that similar to $\epsilon z$-greedy, $\epsilon t$-greedy has polynomial sample complexity in related parameters of the MDP. Our second contribution is a new goal-conditioned dual replay buffer (GDRB), that uses two replay buffers along with an adaptive sampling strategy to differentiate goal-reached and goal-not-reached experience data. These two buffers differ in retention policy, size, and the transitions they store. Our third enhancement is to replace the one-step update in DDPG with the longest $n$-step return for all transitions in an episode. Figure 1 illustrates the innovations of ETGL-DDPG. In Section 4, we evaluate the performance of ETGL-DDPG through extensive experiments on 2D and 3D continuous control benchmarks. As the proposed strategies are orthogonal, we show that each of the three strategies individually improves the performance of DDPG via a thorough ablation study. Furthermore, ETGL-DDPG outperforms current state-of-the-art methods across all tested environments.
@@ 3.1 $\epsilon t$-Greedy: Exploration With Search
To execute the search process, the agent requires access to the environment’s transition function $\mathcal{T}$ of the corresponding MDP. This is used to generate new nodes within the search tree. However, since our exploration strategy is built on DDPG, the model-free algorithm, the transition function $\mathcal{T}$ is not known. Instead, the agent utilizes its replay buffer to advance the search. We briefly discuss the impact of having access to $\mathcal{T}$ on the exploration process in Supplementary Material A.4. We also assume that the agent has a SimHash function $\phi$, which discretizes the large continuous environment. For each state $s$, $n(\phi(s))$ serves as an estimate of the number of visits to a neighbourhood of $s$ throughout the entire learning process.
@@
A policy $\pi$ is called $\epsilon$-optimal if it satisfies $V^{\pi^{*}}(s)-V^{\pi}(s)\leq\epsilon$, for all $s\in S$, where $\epsilon>0$.
@@
For simplicity, when we say an algorithm $\mathcal{A}$ has polynomial sample complexity, we imply that $\mathcal{A}$ is PAC-MDP. The work by Liu & Brunskill (2018) establishes polynomial sample complexity for a uniformly random exploration by bounding the covering length defined in Definition 1. Using this, and considering a limited tree budget $N$, we show that $\epsilon t$-greedy is PAC-MDP. Let’s denote the search tree by $\mathcal{X}$, and the distribution over the generated options in $\mathcal{X}$ as $\mathcal{P}_{{\mathcal{X}}}$. The following Theorem provides a lower bound on option sampling in tree $\mathcal{X}$ under certain condition.
@@ 3.3 Using Longest $n$-step Return
$\small Q(s_{t},a_{t})=\begin{cases}\sum^{T-t}_{k=0}\gamma^{k}r_{t+k},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{$s_{T}$ isa goal state}\\ \\ \sum^{T-t-1}_{k=0}\gamma^{k}r_{t+k}+\gamma^{T-t}Q(s_{T},a_{T}),\ \ \ \ \ \text{otherwise}\end{cases}$ (6)
@@ 4.2 Environment Coverage through Exploration
We now examine how effective $\epsilon t$-greedy is in covering the environment. To do so, we discretize the navigation environments into small cells. A cell is considered visited if the agent encounters a sufficient number of distinct states within it, and the overall environment coverage is quantified as the fraction of visited cells. Figure 4 presents a comparison of environment coverage across different exploration strategies. All strategies except DOIE, which uses Radial Basis Function Deep Q-Network (RBFDQN) (Asadi et al., 2021), use DDPG as their underlying algorithm. RBFDQN is an enhanced DQN variant that incorporates Radial Basis Functions (RBF) to achieve more accurate Q-value approximations in continuous environments. In Wall-maze, $\epsilon t$-greedy achieves 80% coverage, whereas DOIE reaches 60%. $\epsilon z$-greedy covers approximately half of the environment, whereas the remaining methods manage to explore only around 30%. In U-maze, all strategies are successful, covering 80% or more of the environment. Even so, $\epsilon z$-greedy reaches full coverage faster than other methods. In Point-push, none of the methods can fully cover the environment and ultimately achieve nearly the same coverage. The tree budget $N$ serves as an upper bound on the option length in $\epsilon t$-greedy, analogous to its role in $\epsilon z$-greedy when using a uniform distribution $z(n)=\mathbbm{1}_{n\leq N}/N$. We evaluate environment coverage under varying values of $N$ in Section A.7, where results indicate that $\epsilon t$-greedy achieves improved coverage as $N$ increases. In contrast, $\epsilon z$-greedy does not exhibit a consistent improvement in coverage and, in some cases, experiences a decline. This highlights the advantages of directed exploration over undirected approaches. Finally, we present the distribution of final states reached in the episodes to illustrate the order in which the agent visits different regions of the environment (see Appendix A.8).
In $\epsilon z$-greedy, various distributions can be used to select options based on their length. Although a uniform distribution is a straightforward choice, alternative distributions such as the zeta distribution $z(n)\propto n^{-\mu}$ (where $\mu$ controls the decay rate) can also be employed. Empirical results by (Dabney et al., 2020) show that the zeta distribution slightly outperforms the uniform distribution, and moreover, produces a pattern akin to Lévy flights observed in certain ecological foraging models (Baronchelli & Radicchi, 2013). Figure 5 shows the distribution of option lengths generated by $\epsilon t$-greedy during training across all environments, revealing two main observations: (1) moderate-length options have the highest probability of being selected, and (2) probabilities decay as option length increases. While this decay pattern resembles that of the zeta distribution (where length 1 is most likely), $\epsilon t$-greedy instead favors moderate lengths. Note that Although options longer than 12 do occur, their frequency is extremely low and thus not visible in Figure 5.
(a) Wall-maze
(b) U-maze
(c) Point-push
(d) soccer
(e) window-open
(f) press-button
Figure 5: The distribution of options chosen in training. The x-axis represents the length of the options and the y-axis indicates the probability of each length, calculated based on how often each length is chosen across all options.
(a) Wall-maze
(b) U-maze
(c) Point-push
Figure 6: The ablation study for two-component combinations in navigation environments. The success rates are averaged over 5 runs with different random seeds and the shaded areas represent one standard deviation.
@@ 4.3 Effectiveness of Each New Component in ETGL-DDPG
We evaluated the performance of ETGL-DDPG, and now we assess the impact of each component on DDPG separately. Figure 7 presents the results for all environments. $\epsilon t$-greedy demonstrates the most improvement across all environments and is the only method that enhances the performance of DDPG in the Wall-maze, highlighting the critical role of our exploration strategy. GDRB shows a positive impact on DDPG performance in all environments, except for soccer, where DDPG alone outperforms all baselines. Additionally, we replaced reservoir sampling with FIFO as the retention policy in GDRB and observed similar results. The longest n-step return has a positive effect only in U-maze and press-button tasks, while it negatively impacts performance in soccer and Point-push. We attribute this to the inherently high variance of multi-step updates. A comparison of Figures 3 and 7 across all environments shows that ETGL-DDPG consistently outperforms the use of each component individually, supporting the effectiveness of their combination.
To conclude the analysis of each component, we conduct three additional experiments in this section. First, we perform an ablation study by considering all possible two-component combinations, where the omitted component in each combination is replaced with its DDPG counterpart. The results, shown in Figure 6 for the navigation environments, indicate that all combinations outperform DDPG in every environment except Wall-maze, where the agent fails to achieve a non-zero success rate without $\epsilon$-greedy exploration. Moreover, in Point-push, omitting the longest $n$-step return while following the original DDPG strategy causes a noticeable decline in success rates during the later stages of training.
Second, we compare the longest $n$-step return with the method proposed by Meng et al. (2021), who experimented with $n$ ranging from 1 to 8 and tested both the minimum of the $n$-step returns and their average. In their experiments, MMDDPG, which computes the average return over 1 to 8 steps (avg8-step), outperforms other variants in robotic tasks. Here, we replace the longest $n$-step return in ETGL-DDPG with the avg8-step return to evaluate its effectiveness in reward propagation. The results, presented in Appendix A.2, show that both methods perform similarly across most environments, except in Wall-maze, where the longest $n$-step return outperforms the avg8-step approach. Notably, the avg8-step method converges more rapidly, likely due to the stabilizing effect of averaging multiple updates.
Lastly, we compare GDRB with Hindsight Experience Replay (HER) Andrychowicz et al. (2017), a method for sparse-reward tasks that increases feedback to the agent by treating certain states in unsuccessful episodes as imaginary goals. By assigning rewards to these artificially generated goals, HER can leverage unsuccessful episodes more effectively. Figure 8 in Appendix A.3 compares the impact of GDRB and HER on DDPG in navigation environments. In Wall-maze, HER enables the agent to learn from previously unsuccessful trajectories, allowing it to achieve success rates above zero. In the other two environments, both GDRB and HER exhibit similar performance, although HER converges more quickly due to its reward-reshaping mechanism, which provides stronger guidance in the initial stages of training.
@@ 5 Related Work
Intrinsic Motivation. Intrinsic motivation methods  (Burda et al., 2018; Pathak et al., 2017; Ostrovski et al., 2017; Tang et al., 2017) provide a reward bonus for unexplored areas of the state space, enabling the agent to receive more feedback in sparse-reward scenarios. However, since a transition can receive varying rewards at different time steps, these methods make the reward function non-stationary, violating the Markov assumption of MDPs. They are also sensitive to hyperparameters, requiring extensive tuning for each environment (Schäfer et al., 2021). Decoupled RL algorithms (Schäfer et al., 2021; Badia et al., 2019) address this issue by training separate policies for exploration and exploitation, where the exploitation policy optimizes sparse rewards while leveraging data collected by the exploration policy. Although this resolves the non-stationarity issue, it doubles the computational cost. $\epsilon t$-greedy is similar to Tang et al. (2017) in identifying less-explored states using the SimHash function. However, unlike intrinsic motivation methods, $\epsilon t$-greedy does not modify the reward function. Instead, it directs the agent toward less-explored states, preserving the stationarity of the reward function while avoiding the need for additional computation.
@@
Other Exploration Techniques. There are also other strategies to improve exploration without using reward bonuses. Colas et al. (2018) use a policy search process to generate diverse data to improve training of DDPG. Liu et al. (2018) introduce a competition-based exploration method where two agents (A and B) compete with each other. Agent A is penalized for visiting states visited by B, while B is rewarded for visiting states discovered by A. Plappert et al. (2018) directly inject noise into the policy’s parameter space instead of the action space. Eysenbach et al. (2019) build a graph using states in the replay buffer, allowing the agent to navigate distant regions of the environment by applying Dijkstra’s algorithm. $\epsilon t$-greedy shares similarities with Eysenbach et al. (2019) in constructing a tree structure within the state space. However, while Eysenbach et al. (2019) builds a single global graph to connect the start state to the goal state, $\epsilon t$-greedy generates local trees dynamically, focusing on accessing less-explored areas. Ecoffet et al. (2021) introduce Go-Explore which separates exploration and policy optimization into two distinct phases. During the exploration phase, the algorithm systematically explores the environment by maintaining an archive of visited states along with their corresponding trajectories. In the robustification phase, these trajectories are leveraged to train a robust policy using standard reinforcement learning methods, such as PPO. However, a key limitation of Go-Explore is its reliance on resetting trajectories to return to previously visited states, a feature that is often unavailable in many real-world scenarios. Lobel et al. (2022) present Deep Optimistic Initialization for Exploration (DOIE), which improves exploration in continuous control tasks by maintaining optimism in state-action value estimates. Lobel et al. (2023) demonstrate that DOIE can estimate visit counts by averaging samples from the Rademacher distribution instead of using density models. Dey et al. (2024) present COIN, a continual optimistic initialization strategy that extends DOIE to stochastic and non-stationary environments.
@@
Experience Replay Buffer. Rather than uniformly sampling from the buffer, Prioritized Experience Replay (PER) (Schaul et al., 2015) prioritizes transitions in the buffer based on reward, recency, or TD error at the expense of $O(\log N)$ per sample, where $N$ is the buffer size. CER (Zhang & Sutton, 2017) includes the last transition from the buffer to each sampled batch with $O(1)$ complexity. Zhang et al. (2022) learn a conservative value regularizer only from the observed transitions in the replay buffer to improve the sample efficiency of DQN. Pan et al. (2022) theoretically show why PER has a better convergence rate than uniform sampling policy when minimizing mean squared error. Furthermore, Pan et al. (2022) identify two limitations of PER: outdated priorities and insufficient coverage of the state space. Kompella et al. (2023) propose Stratified Sampling from Event Tables (SSET), a method that partitions the experience replay buffer into distinct Event Tables. Each Event Table captures significant subsequences of optimal behavior, allowing for more targeted sampling during training. GDRB aligns conceptually with SSET if achieving a goal is defined as the sole event. However, SSET employs FIFO for inserting new data into its buffers, whereas GDRB uses reservoir sampling for its success-event buffer. Moreover, in SSET, the sampling ratios between buffers are fixed, while GDRB dynamically adjusts these ratios over time to favor the smaller buffer.
Reward Propagation. Reward shaping (Laud, 2004; Hu et al., 2020) creates artificial intermediate rewards to facilitate reward propagation. However, designing appropriate intermediate rewards is hard and often problem-specific. Trott et al. (2019) address this issue by introducing  self-balancing reward shaping in the context of on-policy learning. Their approach defines the local optimum as the closest visited state to the goal and calculates the reward for each state based on its distance from the local optimum and the goal. By dynamically adjusting the local optimum, the agent is guided toward the optimal behavior. To extract more information from an unsuccessful episode, Andrychowicz et al. (2017) introduce imaginary goals. An imaginary goal for state $s$ is a state that is encountered later in the episode. Learning to achieve these imaginary goals helps the agent understand the structure of the environment. Memarian et al. propose Self-Supervised Online Reward Shaping (SORS), a method that alternates between ranking trajectories using sparse rewards and training a classifier to infer a dense reward function from these rankings. The inferred dense rewards are then utilized to update the agent’s policy. Devidze et al. (2024) present a reward informativeness criterion that adaptively constructs interpretable reward functions based on the agent’s current policy.